OpenStudy (anonymous):

*2.3 meters is how many centimeters? **5 1/2 pounds is how many ounces? *** 3 1/4% of 6500 **** 9 to the power of 3

OpenStudy (anonymous):

There are 100 cm in 1 meter. Hence 1 m = 100 cm 2 m = 200 cm 0.1 m = 10 cm 0.3 m = 30 cm 1.3 m = 130 cm Therefore 2.3 m = ..... cm. How many?

OpenStudy (anonymous):

230? Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yay! And my others?

OpenStudy (anonymous):

How many ounces in a pound?

OpenStudy (anonymous):

16... so 44 ounces?

OpenStudy (anonymous):

No. 16 x ( 5 1/2) = 16 x 11/2 = ....

OpenStudy (anonymous):

Now x% of 6500 is \[{x \over 100} 6500\] So 3.5% of 6500 is ....

OpenStudy (anonymous):

9^3 should be easy because \[a^3 = a \times a \times a\] Thus \[9^3 = 9 \times 9 \times 9 = .... what?\]

OpenStudy (anonymous):

227.50000000000003

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can that possibly be right? How on earth did we get so many decimal places.

OpenStudy (anonymous):

84 ounces! I figured out where I went wrong

OpenStudy (anonymous):

So the question on pounds: 16 x ( 5 1/2) = 16 x 11/2 = 8 x 11 (canceling 2) = 88

OpenStudy (anonymous):

5.25 pounds = 84 ounces right?

OpenStudy (anonymous):

**5 1/2 pounds is how many ounces? is 88 but 5.25 pounds is 84 ounces, yes.

OpenStudy (anonymous):

Yay, thanks.